# Reference Coverage Analysis of OpenAlex compared to Web of Science and Scopus

### 29 Jan 2024 v1

@@ 3.1 Source and Total Reference Coverage
Reported Average Source Reference Count 16.867 18.692 7.572
@@
Reported Average Source Reference Count 22.442 20.230 8.134
@@
   Reported Average Source Reference Count 33.416 33.363 34.863
@@
Table 3: Comparison of the reference coverage available in each database, including both the reference coverage numbers from the database providers, and our computed numbers
@@
The internal coverage of OpenAlex cannot be computed for Table 3 as it does not contain a count of total references, however we can infer the internal coverage in the Shared Corpus by assuming either Scopus or WoS contain a definitive reference count. In which case the internal coverage for OpenAlex would be 83.2% when compared to WoS, or 83.6% when compared to Scopus, notably these values lie between those of WoS and Scopus. We cannot perform the same analysis for the other restrictions of the database given the differing database sizes.
When comparing the average total reference counts between WoS and Scopus in Table 3, it initially appears that Scopus outperforms WoS, however when considering records marked as articles they perform more comparably. This trend continues when observing the Shared Corpus and the Shared Corpus with references from 1996 to 2022. Notably here the difference between the source reference count and total reference count decreases as the restrictions are added.
3.2 Discrepancies in Reported Reference Counts
@@
When comparing the reported, pre-calculated average total and source reference counts to the self-calculated ratio of references to publications, we came across a discrepancy in Scopus and OpenAlex. In case of Scopus, reference counts reported by the provider do not always correspond to the actual references in the database. In case of OpenAlex, some references refer to items that do not exist in OpenAlex, i.e. are deleted. The latter references are not included in the pre-calculated values. The discrepancies between both types of calculation can be seen in Table 4.
@@
For Scopus and OpenAlex, we selected the publications where either the pre-calculated total ‘reference count’ in Scopus and ‘source reference count’ in OpenAlex are not equal to the respective number of entries in the databases’ reference table. We computed the average of the reported/pre-calculated counts, and then compared this to the ratio of references to publications while excluding in both cases publications where pre-calculated and self-calculated total reference count (in Scopus) or source reference count (in OpenAlex) do not match. Once this has been done, the resulting averages then only differ at the 12th to 14th decimal place. We therefore conclude that for both database discrepancies between pre- and self-calculated reference counts arise due to erroneous data.
@@
This result raises issues with both proprietary and open source bibliometric databases and should be considered when utilising these databases for citation analysis, and likely merits a deeper analysis.
@@ 3.3 Metadata Coverage
Figure 2 highlights the metadata coverage analysis results between OpenAlex and the two proprietary databases, WoS and Scopus within the Shared Corpus. The x-axis represents OpenAlex, while the y-axis corresponds to WoS (left) and Scopus (right). The points represent the percentage coverage of the relevant indicator per journal.
@@ 3.4 Discussion
This study demonstrates the source reference coverage in OpenAlex to be comparable to that in WoS and Scopus for the records from 2015-2022 which lie in the intersection of all three databases, both in general and when restricting to references from 1996 onwards. Table 3 also demonstrates that within the Shared Corpus, there are on average 6.4 to 6.2 references captured in the total reference count by WoS and Scopus (respectively) that OpenAlex does not capture in its source reference count.
We note that the reported figures for average total reference count in Scopus and average source reference count in OpenAlex do not equal the average number of references in each database. Similarly we note that matching by DOI in these databases is not a strictly one-to-one match, with some DOIs referring to multiple records in all three databases (cf. [16]). This observation indicates that there may be errors in naïve computation of the reference counts in each database and we believe it merits further study and caution when replicating these computations.
Although metadata coverage relating to abstract information is lower than in WoS and Scopus, the share of records with abstracts in OpenAlex is nevertheless higher than in Crossref as stated by Kramer in [17]. Kramer also notes that, at the time of writing, the large publishers Elsevier, Taylor & Francis and IEEE did not openly share abstracts via Crossref. But OpenAlex also acknowledged legal issues, which resulted in the representation of abstracts as inverted index as well as in the removal of some abstracts. [^5] https://groups.google.com/g/openalex-users/c/ptFDD7qWvYw/m/kXWDG3o5BAAJ
@@ 4 Limitations and Outlook
We also do not check the accuracy with which the databases match references to publications, which can be seen as the prerequisite for the internal coverage indicator we use. Some studies analyse the accuracy of the database matching algorithms either on the basis of manual sample evaluations and/or in comparison with their own algorithms (e.g. [18].)
@@
Acknowledgements
This work was funded by the Federal Ministry of Education and Research via funding numbers: 16WIK2301B / 16WIK2301E, The OpenBib project. We acknowledge support by Federal Ministry of Education and Research, Germany under grant number 01PQ17001, the Competence Network for Bibliometrics.

### 26 Feb 2024 v2

@@ 3.1 Total and Source Reference Coverage
Pre-calculated Average Source Reference Count 16.867 18.692 7.572
@@
Pre-calculated Average Source Reference Count 22.442 20.230 8.134
@@
   Pre-calculated Average Source Reference Count 33.416 33.363 34.863
@@
Table 3: Comparison of the reference coverage available in each database, including the reference counts from the database providers, and our computed counts
@@
The internal coverage of OpenAlex cannot be computed for Table 3 as it does not contain all references, respectively a total reference count. However, we can infer OpenAlex’ internal coverage in the Shared Corpus by assuming either Scopus or WoS contain a definitive reference count. In this case, the internal coverage for the last segment (comprising the 1996-2022 restriction to reference publication years) for OpenAlex would be 83.2% when related to WoS’ total reference count, or 83.6% when related to Scopus’ reference count, notably these values lie between those of WoS and Scopus. We cannot perform the same analysis on all comparisons given the differing database sizes.
3.2 Discrepancies between Reference Counts and Reference Data
@@
When comparing the reported and pre-calculated average total and source reference counts to an alternatively self-calculated ratio of all references to all publications, we came across discrepancies in Scopus and OpenAlex. In case of Scopus, reference counts reported by the provider do not always correspond to the actual references in the database, a phenomenon confirmed by Elsevier in informal communication as being caused by inconsistent supplier ingestions. In case of OpenAlex, some references refer to items that do not exist in OpenAlex, i.e. are deleted. The latter references are not included in the pre-calculated values. The discrepancies between both types of calculation can be seen in Table 4.
@@
For further verification, we selected the publications in Scopus and OpenAlex where either the pre-calculated total ‘reference count’ in Scopus and ‘source reference count’ in OpenAlex were not equal to the respective number of entries in the databases’ reference table. We then computed the averages of the reported/pre-calculated counts, and compared this to the ratio of references to publications while excluding in both cases the identified publications where reference count (in Scopus) or source reference count (in OpenAlex) do not correspond to the actual number of references. Once this has been done, the resulting averages then only differ at the 12th to 14th decimal place. We therefore conclude that for both databases discrepancies between reference counts and actual reference numbers are due to erroneous data. While in OpenAlex both our pre- and self-calculated source reference counts are consistent to our concept as we only count as source references those whose target items are actually in the database, the situation is more complicated in the case of Scopus: In our averages, we first use the reference counts supplied by the provider, which do not always match (but are probably more correct than) the references actually supplied, while in the last segment, where we calculate the count ourselves with references restricted to the 1996-2022 time window, we can only do this on the basis of the references actually supplied.
@@
The detected discrepancies raise issues with both proprietary and open source bibliometric databases and should be considered when working with the references, and likely merits a deeper analysis.
@@ 3.3 Metadata Coverage
Figure 2 highlights the metadata coverage analysis results between OpenAlex and the two proprietary databases, WoS and Scopus, within the Shared Corpus. The x-axis represents OpenAlex, while the y-axis corresponds to WoS (left) and Scopus (right). The points represent the percentage coverage of the relevant indicator per journal.
@@ 3.4 Discussion
This study demonstrates the source reference coverage in OpenAlex to be comparable to that in WoS and Scopus for comparatively newer records which lie in the intersection of all three databases, both in general and when restricting to references from 1996 onwards. On the one hand, this could be seen as an indicator of good quality bibliometric core data. On the other hand, OpenAlex does not have the highest internal coverage, although it is by far the largest database, so it would actually be plausible that higher proportions of the referenced publications are themselves part of the database. In this respect, the Scopus coverage policy seems to be a bit more effective. However, one possible factor could also be that a comparatively poorer reference-matching algorithm misses a noticeable amount of actual source references. From table 3 it can be inferred that within the Shared Corpus, there are on average 6.4 to 6.2 references captured in the total reference count by WoS and Scopus (respectively) that OpenAlex does not capture in its source reference count. The fact that OpenAlex does not systematically include non-source references, as well as complete reference strings, limits the flexibility of using and exploring the data source: It does not allow researchers or bibliometric centers to apply their own reference matching algorithms or to analyse non-source references as such.
The study also revealed data errors in Scopus and OpenAlex. The reported figures for reference counts in Scopus do not correspond to the actual numbers of references in the database, and OpenAlex is inconsistent in its handling of references as it does not systematically comprise all non-source references, but references to deleted source items. Similarly, we note that all databases, to a different degree, comprise cases where DOIs refer to multiple records–cf. [5]. We believe it merits further study and caution when replicating these computations.
Although metadata coverage relating to abstract information is lower than in WoS and Scopus, the share of records with abstracts in OpenAlex is nevertheless higher than in Crossref as stated in [9]. Kramer also notes that, at the time of writing, the large publishers Elsevier, Taylor & Francis and IEEE did not openly share abstracts via Crossref. But OpenAlex also acknowledged legal issues, which resulted in the representation of abstracts as inverted index as well as in the removal of some abstracts. [^5] https://groups.google.com/g/openalex-users/c/ptFDD7qWvYw/m/kXWDG3o5BAAJ
@@ 4 Limitations and Outlook
We also do not check the accuracy with which the databases match references to publications, which can be seen as the prerequisite for the internal coverage indicator we use. Some studies analyse the accuracy of the database matching algorithms either on the basis of manual sample evaluations and/or in comparison with their own algorithms for example, in [13].
@@
Declarations
[…]
Funding This work was funded by the Federal Ministry of Education and Research via funding numbers: 16WIK2301B / 16WIK2301E, The OpenBib project. We acknowledge support by Federal Ministry of Education and Research, Germany under grant number 01PQ17001, the Competence Network for Bibliometrics.
@@
[…]
Conflicts of interests/Competing interests
Philipp Mayr, the co-author of this paper, has a conflict of interest because he serves on the editorial board of the journal Scientometrics.